OpenStudy (anonymous):

Help with these 3 questions Pleasee! I can't fail this test! 1. x - 2/7 =3 A. x = -3 2/7 B. x = -2 5/7 C. x= 2 5/7 D. x= 3 2/7 2. 1/4 + w =6 1/3 A. w = 6 1/2 B. w = 6 1/7 C. w = 6 2/7 D. w = 6 7/12 3. -3 2/5 + B = 8 1/5 A. b = 11 3/5 B. b= 4 4/5 C. b= 5 3/5 D. b= 11 1/5 Thanks!!! :)

OpenStudy (anonymous):

:l

OpenStudy (texaschic101):

x - 2/7 = 3 x = 2/7 + 3 x = 2/7 + 21/7 x = 23/7 x = 3 2/7 (D) ============== 1/4 + w = 6 1/3 w = 19/3 - 1/4 w = 76/12 - 3/12 w = 73/12 w = 6 1/12 ---> I am getting a different answer then your choices ============== -3 2/5 + b = 8 1/5 - 17/5 + b = 41/5 b = 41/5 + 17/5 b = 58/5 b = 11 3/5 (A)

OpenStudy (anonymous):

Thanks a ton!!!

OpenStudy (texaschic101):

your welcome...a ton...lol :)